OpenStudy (anonymous):

graph the linear equation y=1/3x-4

OpenStudy (anonymous):

i meant graph the linear function is that the same thing

OpenStudy (anonymous):

y = 1/3x - 4 Multiply on both sides by 3 3y = x - 12 x-3y-12=0 Assume x=0, then on solving y=-4 one point (0,-4) Assume y=0, by solving, x=12 another point (12,0) 

OpenStudy (anonymous):

thank you can you help me with another one

OpenStudy (anonymous):

y=-2x+3

OpenStudy (anonymous):

y=-2x+3 2x+y-3=0 Assume x=0, then y=3 point(0,3) Assume y=1, then x=1 point(1,1) 

OpenStudy (anonymous):

thank you I need to pass math to pass my grade and I don't understand any of this stuff I have some questions can you still help me

OpenStudy (anonymous):

oh sure. but you have to learn how to do all these.

OpenStudy (anonymous):

i will i am just really stressed out right now cant think straight

OpenStudy (anonymous):

a 154lb person burns 420 calories per hour riding an exercise bike at a rate 15mi/h. Write a function rule to represent the total calories burned over time by that person. Explain how the information in the problem relates to the function.

OpenStudy (anonymous):

explain how to write a function rule from the table below. then write a function rule . x|2|4|6 y|1|0|-1

OpenStudy (anonymous):

last question please help me

OpenStudy (anonymous):

never mind got the last question. It is actually one I understand

OpenStudy (anonymous):

hmm... fine...